# Spatial structure facilitates evolutionary rescue by cost-free drug resistance

### v1

@@ 3.5 Dynamics at the single-deme level
In Figure 3, we analyzed the time evolution of mean and variance across replicates of the total number of mutants in the population. In Figure S4, we analyze these quantities at the single-deme level. Since the clique is a symmetric structure such that all demes are equivalent, we select deme 1 in each of our simulation replicates for this analysis. Figure S4A shows the average number of mutants in this deme as a function of time. This quantity is the same in the structures with different migration rates, as is the case at the population level (see Figure 3A). Conversely, Figure S4B shows that the variance of the number of mutants in one deme differs from its population-level counterpart (see Figure 3B). Indeed, at the deme level, there is no visible difference between different population structures as far as the variance across replicates of the number of mutants in a deme is concerned, see Figure S4B. Thus, the local mutant dynamics are the same for all migration rates. It is the global dynamics at the population level, combining all demes, which gives rise to the impact of spatial structure that we evidenced in this work.
@@
Figure S4: Mean and variance of the number of mutants at the single-deme level. Panel A: the mean number of mutants in deme 1 is shown as a function of time in the absence of drug, for different values of the migration rate. Panel B: the variance across simulation replicates of the number of mutants in deme 1 is shown as a function of time in the absence of drug, for different values of the migration rate. Data is obtained from $10^{4}$ simulation replicates in each case. The case of an isolated deme, i.e. of a well mixed population with carrying capacity $K$ (“WMP ($K=100$)”), is shown for reference. Parameter values (in both panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
4 Lattice and star structures with sensitive inoculum
@@
Our work mainly focuses on a minimal model of spatially structured populations where all demes are equivalent and connected to one another by identical migration rates. This structure is known as the island model, the clique or the fully connected graph. Let us now extend our study to different graph structures. We first consider the square lattice (see Figure S5A). It is a symmetric structure, like the clique, but migrations from each deme are restricted to its four nearest neighbors. More precisely, in the lattice, as in the clique, in each deme the inward migrations balance the outward migrations, i.e. for each $i$ we have $\sum_{j}m_{ij}=\sum_{j}m_{ji}$, where $m_{ij}$ denotes the migration rate from deme $i$ to deme $j$: this means that the lattice is a circulation [lieberman_evolutionary_2005, marrec_toward_2021, abbara_frequent_2023]. The per capita migration rate in the lattice is denoted by $\gamma_{L}$, equal in all directions. We further consider the star (see Figure S5B), which is less symmetric in the sense that the central deme is different from the leaf demes. Migrations from a leaf to the center occur at per capita rate $\gamma_{I}$, while migrations from the center to the leaf occur at per capita rate $\gamma_{O}$. We define the migration asymmetry parameter as $\alpha=\gamma_{I}/\gamma_{O}$. If $\alpha=1$, the star is a circulation, while for all other values of $\alpha$ it is not, which impacts the fixation probability of a mutant in the whole population [marrec_toward_2021, abbara_frequent_2023].
@@
Figure S5: Survival probability of a bacterial population to biostatic drug, for different spatial structures. Panel A: Lattice structure with $D=9$ and migrations $\gamma_{L}$ to each nearest neighbor. Panel B: Star structure with $D=9$ and migration rates $\gamma_{O}$ from the center to a leaf and $\gamma_{I}=\alpha\gamma_{O}$ from a leaf to the center. Panel C: Survival probability of a population starting from a sensitive inoculum versus drug addition time $t_{add}$, for different spatial structures (fully subdivided population with no migrations, “$\gamma=0$”; lattice; clique; star with different migration asymmetries). Parameter values: $D=16$, $K=100$, $f_{R}=1$, $f_{S}=1$ before drug addition, $f_{S}=0$ when antibiotic is added, $g=0.1$, $\mu=10^{-5}$, $\gamma=10^{-6}$ for the clique, $\gamma_{L}=3.75\times 10^{-6}$ for the lattice, $\gamma_{O}=15\times 10^{-6}$ for the star, and $\gamma_{I}=\alpha\gamma_{O}$. Each result is obtained from $10^{3}$ simulation replicates. Error bars show the 95% confidence intervals.
@@
Mutant fixation in the fastest deme is at the root of the impact of population structure on resistance evolution that we evidenced here, in particular in Figure 2. Suppose that a successful mutant appears in the fastest deme: wild-type individuals from connected demes can migrate to the fastest deme and take over there, which takes the population back to the initial fully wild-type state. Thus, to choose migration rates that make different structures comparable, we set migration rates so that the rate of invasion of one deme by the connected demes is the same for two structures. For the lattice, this leads to $\gamma_{L}=\gamma(D-1)/4$, where $\gamma$ is the per capita migration rate in the clique. Note that this condition also yields an equal total exchange rate in the clique and the lattice. For the star, we first remark that a successful mutant is $(D-1)$ times more likely to first appear in a leaf than in the center. Neglecting the case where a successful mutant first appears in the center leads to the condition $\gamma_{O}=(D-1)\gamma$. Note that this condition does not coincide with setting equal overall exchanges between demes in the star and in the clique.
@@
Figure S5C shows the survival probability of bacterial populations with different spatial structures. We observe that the fully subdivided population with $\gamma=0$ has a larger survival probability than all others. This generalizes our results in Figure 2. It can be attributed to the fact that wild type bacteria cannot re-invade the fastest deme in the absence of migrations. Moreover, we find that the lattice and the clique with the same invasion rate do not feature any statistical difference in their survival probabilities. More generally, we observe that the specific graph structure of the population does not significantly affect the survival probability if the addition time $t_{add}$ of antibiotic satisfies $t_{add}\lesssim\langle t_{afF}\rangle$, where $\langle t_{afF}\rangle$ is the average time it takes for a successful mutant to appear in the fastest deme. The minor differences observed between the stars with different asymmetries and between them and other structures for $t_{add}>\langle t_{afF}\rangle$ can be attributed to the fact that more than one deme may then have fixed mutants. Recall indeed that our matching condition for re-invasion was constructed assuming that only one deme was mutant – and that for the star it did not match the condition of same total exchanges between demes.
@@ 5 Clique population structure with mutants in the inoculum
In Figure S6, we show the survival probability of a bacterial population to biostatic antibiotic treatment versus the addition time $t_{add}$ of antibiotic for different migration rates, in the case where each deme is inoculated with 5% mutants. Our simulation results show that the survival probability decreases when $t_{add}$ is increased in this scenario. This was expected, because new mutants do not appear, and the mutants initially present in the inoculum may go extinct. However, the spatial structure of the population still plays a crucial role. Indeed, as the migration rate $\gamma$ decreases, the survival probability increases. This is reminiscent of our result with a sensitive inoculum (see Figure 2). We added the expected probabilities to still have mutants in the system as dashed lines. In a well-mixed population with initial fraction of mutants $x_{ini}$, the probability of having mutants in the long term is equal to their fixation probability, i.e. to $x_{ini}$ since we are considering neutral mutants. In a fully subdivided population with no migrations, this probability is equal to $1-(1-x_{ini})^{D}$, which is the probability that there is at least one deme where mutants fix. In a spatially structured population with nonzero migration rates, the survival probability gradually converges to the well-mixed population one because of migrations.
Figure S6: Survival probability of a structured bacterial population with mutants in the inoculum upon addition of a biostatic drug. The survival probability is plotted versus the addition time $t_{add}$ of antimicrobial. We consider a clique population structure with various migration rates. Results for a well-mixed population with same total size (“WMP”) are presented as reference. Horizontal dashed lines: analytical predictions of the probabilities to still have mutants in the system in the long term (dark blue: well-mixed population; yellow: fully subdivided population with $\gamma=0$). Data is obtained from $10^{4}$ replicate simulations in each case. Parameter values: $K=100$, $D=10$, $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=1$, $g=0.1$, $\mu=0$. Initial percentage of mutants in each deme (and in the well-mixed population): 5%.

### v2

@@ 3.5 Dynamics at the single-deme level
In Fig 3, we analyzed the time evolution of mean and variance across replicates of the total number of neutral mutants in the population. In Fig S4, we analyze these quantities at the single-deme level, always for neutral mutants. Since the clique is a symmetric structure such that all demes are equivalent, we select deme 1 in each of our simulation replicates for this analysis. Fig S4A shows the average number of mutants in this deme as a function of time. This quantity is the same in the structures with different migration rates, as is the case at the population level (see Fig 3A). Conversely, Fig S4B shows that the variance of the number of mutants in one deme differs from its population-level counterpart (see Fig 3B). Indeed, at the deme level, there is no visible difference between different population structures as far as the variance across replicates of the number of mutants in a deme is concerned, see Fig S4B. Thus, the local neutral mutant dynamics are the same for all migration rates. It is the global dynamics at the population level, combining all demes, which gives rise to the impact of spatial structure that we evidenced in this work.
@@
Figure S4: Mean and variance of the number of mutants at the single-deme level. Panel A: the mean number of mutants in deme 1 is shown as a function of time in the absence of drug, for different values of the migration rate. Panel B: the variance across simulation replicates of the number of mutants in deme 1 is shown as a function of time in the absence of drug, for different values of the migration rate. Data is obtained from $10^{4}$ simulation replicates in each case. The case of an isolated deme, i.e. of a well mixed population with carrying capacity $K$ (“WMP ($K=100$)”), is shown for reference. Parameter values (in both panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$ (neutral mutants), $g=0.1$, $\mu=10^{-5}$.
3.6 Population composition versus time for different migration rates
In Fig 3C, we analyzed the composition of the population at the time featuring the largest inter-replicate variance for the well-mixed population, namely $t=5\times 10^{5}$, in spatially structured populations with different migration rates, in the case of neutral R mutants. In Fig S5, we report how this composition evolves in time, in the absence of drug, for spatially structured populations with different migration rates. As in Fig 3C, we focus on the number of mutants in each deme to categorize the population composition.
Figure S5: Population composition versus time for different population structures. Population composition in the absence of drug in the well-mixed population (“WMP”, Panel A), in the fully subdivided population (“SP”, Panel B), and in the clique-structured population with same total size for three values of the migration rate $\gamma$ (Panels C-E). The four categories of population composition reported here are based on the number of mutants per deme. They are the same as in Fig 3C, and are defined in the caption of that figure. The x-axis tick corresponding to $t=5\times 10^{5}$, considered in Fig 3C, is highlighted in purple. Data is obtained from $10^{4}$ simulation replicates in each case. Parameter values (in all panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
Recall that here, the population is composed of wild-type sensitive bacteria and of neutral resistant mutants that arise with probability $\mu$ upon division of wild-types. Fig S5 shows that when the per capita migration rate $\gamma$ is decreased, there is a longer phase where the population has an intermediate composition, with some demes harboring a substantial number of resistant mutants (“Big R pop.” category). This is consistent with our above discussion in Section 3.3 and with our results in Fig S3. The fixation of mutants occurs gradually, deme by deme, in structured populations with small migration rates. Furthermore, substantial mutant numbers exist earlier in these structured populations. We further observe in Fig S5 that a per capita migration rate $\gamma=10^{-3}$ results in an evolution of population composition akin to that of a well-mixed population, a per capita migration rate $\gamma=10^{-7}$ yields an evolution of population composition similar to that of a fully subdivided population with $\gamma=0$. Structures with intermediate values of $\gamma$ feature an intermediate evolution of population composition.
3.7 Colonization timescales after drug is added
In the main text, we discuss the time $\langle t_{\textrm{c mig}}\rangle$ for R mutants to colonize the next empty deme after drug is added, see Eq 7. In Table S1, we report values of $\langle t_{\textrm{c mig}}(k)\rangle$ when $k=1$ or 5 demes are already mutant, for different values of the migration rate $\gamma$, using the same parameter values as in Fig 5. With these parameters, the decay time upon drug addition of a well-mixed population comprising $N^{*}$ S bacteria is $\tau_{S}=50.8$. Recall that the smallest value of $\langle t_{\textrm{c mig}}(k)\rangle$ is obtained when $k=D/2=5$. Thus, the results of Table S1 show that for $\gamma=10^{-6}$, which is the value used in Fig 5, we have $\tau_{S}\ll\langle t_{\textrm{c mig}}(k)\rangle$ for all $k$.
$\gamma=10^{-4}$ $\gamma=10^{-5}$ $\gamma=10^{-6}$ $\gamma=10^{-7}$
$\langle t_{\textrm{c mig}}(k=1)\rangle$ $13.7$ $1.37\times 10^{2}$ $1.37\times 10^{3}$ $1.37\times 10^{4}$
$\langle t_{\textrm{c mig}}(k=5)\rangle$ $4.93$ $49.3$ $4.93\times 10^{2}$ $4.93\times 10^{3}$
Table S1: Numerical evaluation of $\langle t_{\textrm{c mig}}\rangle$ from Eq 7 for $k=1$ and $k=5$. Parameter values: $K=100$, $D=10$, $g=0.1$, $f_{R}=1$ (neutral mutants), as in Fig 5.
The total colonization time $\langle t_{\textrm{c tot}}\rangle$ can be obtained by summing $\langle t_{\textrm{c mig}}\rangle$ over the $D-1$ steps needed to sequentially colonize the structured system. Using Eq 7, we thus obtain:
$\langle t_{\textrm{c tot}}\rangle=\sum_{k=1}^{D-1}\langle t_{\textrm{c mig}}(k)\rangle=\frac{2(\Gamma+\psi(D))}{D\gamma N^{*}(1-g/f_{R})}\,,$ (S36)
where $\Gamma$ is the Euler gamma constant, while $\psi$ is the digamma function.
4 Lattice, star and line structures with sensitive inoculum
Our work mainly focuses on a minimal model of spatially structured populations where all demes are equivalent and connected to one another by identical migration rates. This structure is known as the island model, the clique or the fully connected graph. Let us now extend our study to different graph structures.
Different structures.
@@ 4 Lattice, star and line structures with sensitive inoculum
We first consider the grid or square lattice (see Fig S6A). It is a symmetric structure, like the clique, but migrations from each deme are restricted to its four nearest neighbors, thus only allowing local migrations. In the lattice, as in the clique, in each deme the inward migrations balance the outward migrations, i.e. for each $i$ we have $\sum_{j}\gamma_{ij}=\sum_{j}\gamma_{ji}$, where $\gamma_{ij}$ denotes the migration rate from deme $i$ to deme $j$: this means that the lattice is a circulation [lieberman_evolutionary_2005, marrec_toward_2021, abbara_frequent_2023]. The per capita migration rate in the lattice is denoted by $\gamma_{G}$, equal in all directions.
We further consider the star (see Fig S6B), comprising a central deme connected to $D-1$ leaves [marrec_toward_2021]. All leaves are assumed to be equivalent. It is less symmetric than the clique or the square lattice in the sense that the central deme is different from the leaf demes. Migrations from a leaf to the center occur at per capita rate $\gamma_{I}$, while migrations from the center to the leaf occur at per capita rate $\gamma_{O}$. We define the migration asymmetry parameter as $\alpha=\gamma_{I}/\gamma_{O}$. If $\alpha=1$, the star is a circulation, while for all other values of $\alpha$ it is not, which impacts the fixation probability of a mutant in the whole population [marrec_toward_2021, abbara_frequent_2023].
Finally, we consider a line (see Fig S6C). Migrations from the left to the right occur at per capita rate $\gamma_{R}$, from the right to the left at rate $\gamma_{L}$. An asymmetry parameter is defined for the line as well: $\alpha=\gamma_{R}/\gamma_{L}$. The structure is symmetric under the simultaneous transformations $\alpha\to 1/\alpha$, $\gamma_{R}\to\gamma_{L}$, and $\gamma_{L}\to\gamma_{R}$. As for the star, setting $\alpha=1$ results in a circulation, while $\alpha\neq 1$ impacts the mutant fixation probability [servajean_impact_2024].
@@
Figure S6: Survival probability of a bacterial population to biostatic drug, for different spatial structures. Panels A, B, C: Schematics of the spatial structures considered, shown with $D=9$ demes for visualization ease: a square lattice with migrations $\gamma_{G}$ to each nearest neighbor (A), a star with and migration rates $\gamma_{O}$ from the center to a leaf and $\gamma_{I}=\alpha\gamma_{O}$ from a leaf to the center (B), and a line with migration rates $\gamma_{L}$ to the left and $\gamma_{R}=\alpha\gamma_{L}$ to the right (C). Panels D, E: Survival probability of a population starting from a sensitive inoculum versus drug addition time $t_{\textrm{add}}$, for different spatial structures: line with different migration asymmetries and lattice (D), star with different migration asymmetries (E). Results for a fully subdivided population with no migrations (“$\gamma=0$”), a clique, and a well-mixed population are shown in both panels for reference. For the fully subdivided and well-mixed populations, the prediction from Eq 4 is shown as a solid line. Other lines are guides for the eye. Parameter values: $D=16$, $K=100$, $f_{R}=1$ (no resistance cost), $f_{S}=1$ before drug addition, $f_{S}=0$ when antibiotic is added, $g=0.1$, $\mu=10^{-5}$, $\gamma=10^{-6}$ for the clique, $\gamma_{G}=3.75\times 10^{-6}$ for the lattice, $\gamma_{O}=15\times 10^{-6}$ and $\gamma_{I}=\alpha\gamma_{O}$ for the star, $\gamma_{L}=\gamma_{R}=7.5\times 10^{-6}$ for the line with $\alpha=1$, $\gamma_{L}=1.2\times 10^{-5}$ and $\gamma_{R}=\alpha\gamma_{L}$ for the line with $\alpha=4$. Each result is obtained from $10^{4}$ simulation replicates.
Choosing migration rates.
@@
Mutant fixation in the fastest deme is at the root of the impact of population structure on resistance evolution that we evidenced here, in particular in Fig 2. Suppose that a successful R mutant appears in the fastest deme: S individuals from connected demes can migrate to the fastest deme and take over there, which takes the population back to the initial fully S state. Thus, to make different structures as comparable as possible, we set migration rates so that the rate of invasion of one deme by the connected demes is the same across structures.
For the square lattice or grid, this leads to $\gamma_{G}=\gamma(D-1)/4$, where $\gamma$ is the per capita migration rate in the clique. Note that this condition also yields an equal total exchange rate in the clique and the lattice.
For the star, we first remark that a successful mutant is $(D-1)$ times more likely to first appear in a leaf than in the center. Neglecting the case where a successful mutant first appears in the center leads to the condition $\gamma_{O}=(D-1)\gamma$. Note that this condition does not coincide with setting equal overall exchanges between demes in the star and in the clique.
Finally, for the line, we note that for $D\gg 1$, successful mutants are much more likely to appear in a deme that is not at an end of the line. We thus choose migration rates such that the rate of invasion of one non-end deme by neighboring ones in the line is the same as in the clique: $\gamma_{L}=\gamma(D-1)/(\alpha+1)$, and $\gamma_{R}=\alpha\gamma_{L}$. Note that this condition slightly differs from setting equal overall exchanges in the line and in the clique, which would give $D$ instead of $D-1$ in the numerator of the expression of $\gamma_{L}$.
Simulation results.
@@
Fig S6C shows the survival probability of bacterial populations with different spatial structures with $D=16$ demes. We observe that the fully subdivided population with $\gamma=0$ has a larger survival probability than all others. This generalizes our results in Fig 2. It can be attributed to the fact that S bacteria cannot re-invade the demes that fixed resistance in the absence of migrations. Moreover, we do not observe statistically significant differences between survival probabilities in the lattice and the clique with the same invasion rate. More generally, we observe that the specific graph structure of the population does not significantly affect the survival probability if the addition time $t_{\textrm{add}}$ of antibiotic satisfies $t_{\textrm{add}}\lesssim\langle t_{afF}(N^{*},D)\rangle$, where $\langle t_{afF}(N^{*},D)\rangle$ is the average time it takes for a successful mutant to appear in the fastest deme. The minor differences observed between the stars with different asymmetries, between them and other structures, and between lines and other structures for $t_{\textrm{add}}>\langle t_{afF}(N^{*},D)\rangle$, can be attributed to the imperfections of our matching condition for the line and for the star, and to the fact that more than one deme may then have fixed mutants. Recall indeed that incoming migrations to the center of the star and to the end demes of the line were not matched, that our matching condition for re-invasion was constructed assuming that only one deme was mutant, and that for the star and the line, our condition did not match the condition of same total exchanges between demes.
@@ 5 Clique population structure with mutants in the inoculum
In Fig S7, we show the probability that a bacterial population survives biostatic antibiotic treatment versus the addition time $t_{\textrm{add}}$ of antibiotic for different migration rates, in the case where each deme is inoculated with 5% mutants. Our simulation results show that the survival probability decreases when $t_{\textrm{add}}$ is increased in this scenario. This was expected, because new mutants do not appear, and the mutants initially present in the inoculum may go extinct. However, the spatial structure of the population still plays a crucial role. Indeed, as the migration rate $\gamma$ decreases, the survival probability increases. This is reminiscent of our result with a sensitive inoculum (see Fig 2). We added the expected probabilities to still have mutants in the system as dashed lines. In a well-mixed population with initial fraction of mutants $x_{\textrm{ini}}$, the probability of having mutants in the long term is equal to their fixation probability, i.e. to $x_{\textrm{ini}}$ since we are considering neutral mutants. In a fully subdivided population with no migrations, this probability is equal to $1-(1-x_{\textrm{ini}})^{D}$, which is the probability that there is at least one deme where mutants fix. In a spatially structured population with nonzero migration rates, the survival probability gradually converges to the well-mixed population one because of migrations.
Figure S7: Survival probability of a structured bacterial population with mutants in the inoculum upon addition of a biostatic drug. The survival probability is plotted versus the addition time $t_{\textrm{add}}$ of antimicrobial. We consider a clique population structure with various migration rates. Results for a well-mixed population with same total size (“WMP”) are presented as reference. Horizontal dashed lines: analytical predictions of the probabilities to still have mutants in the system in the long term (dark blue: well-mixed population; yellow: fully subdivided population with $\gamma=0$). Data is obtained from $10^{4}$ replicate simulations in each case. Parameter values: $K=100$, $D=10$, $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=1$, $g=0.1$, $\mu=0$. Initial percentage of mutants in each deme (and in the well-mixed population): 5%.
6 Concrete examples of spatially structured populations
In hosts, different organs can feature bacterial colonization and possess spatial structures with numerous demes. Here, we provide some details on the orders of magnitudes mentioned in the Discussion.
Estimate of the number of intestinal crypts.
[…]
Estimate of the number of skin pores.
[…]